OpenStudy (anonymous):

Check my work? Sophia currently has an account balance of $1,347.11. Her initial deposit on the account was $843 and it earned 2.3% simple interest. How long has Sophia held the account? 16 years 26 years 43 years 69 years

OpenStudy (anonymous):

1,347.11 - 843 = 504.11 2.3% of 843 =19.389 504.11/19.389 = 25.999, round it up to 26

OpenStudy (cj49):

u did fine mate

OpenStudy (anonymous):

Alright thanks :)

ganeshie8 (ganeshie8):

843*0.023*26 = 504.11 843 + 504.11 = 1347.11

OpenStudy (anonymous):

Thanks for the check gane

ganeshie8 (ganeshie8):

np :)